OpenStudy (anonymous):

((e^(x)-e^(-1))^2)/e^(2x) write expression so only one exponential function appears

hartnn (hartnn):

Lets first simplify numerator only can u expand ((e^(x)-e^(-1))^2) using (a+b)^2 = a^2+2ab+b^2 formula? what do u get ?

OpenStudy (anonymous):

(e^x^2)-2+(e^(-x)^2)

OpenStudy (anonymous):

sorry, the expression is ((e^(x)-e^(-x))^2)/e^(2x)

hartnn (hartnn):

ok, then thats correct expansion of numerator. but u can write \((e^x)^2=e^{2x}\) ok ? so your numerator will be \(\huge e^{2x}-2+e^{-2x}\) got this ?

OpenStudy (anonymous):

ok

hartnn (hartnn):

now u can distribute denominator's e^2x to each term of numerator . \(\frac{e^{2x}}{e^{2x}}-\frac{2}{e^{2x}}+\frac{e^{-2x}}{e^{2x}}\) ok? and use \(\huge \frac{e^{-2x}}{e^{2x}}=e^{-2x-2x}=?\) so what u get ?

OpenStudy (anonymous):

\[e ^{-4x}\]

hartnn (hartnn):

so your expression will be \(\huge 1-2e^{-2x}+e^{-4x} \) got this?

OpenStudy (anonymous):

ok

hartnn (hartnn):

now again using (a+b)^2 = a^2+2ab+b^2 formula., taking a=1 and \(b= e^{-2x}\) we can simplify that as \(\huge (1-e^{-2x})^2\) any doubts for this?

OpenStudy (anonymous):

No

hartnn (hartnn):

then thats your final answer.

OpenStudy (anonymous):

Thank you!

hartnn (hartnn):

welcome :)